OpenStudy (anonymous):

sqaure root 20 + square root 40

OpenStudy (anonymous):

2 radical 5 + 2 radical 10

OpenStudy (anonymous):

but then simplify it to what ? cuz thats not one of my anwsers

OpenStudy (anonymous):

hmm what are your answers

OpenStudy (anonymous):

sr 10 , 5sr5 , 6 sr 5 or cant combine terms

OpenStudy (radar):

\[\sqrt{4}\sqrt{5}+\sqrt{4}\sqrt{10}=2\sqrt{5}+2\sqrt{5}\sqrt{2}=\sqrt{5}(2+\sqrt{2)}\]

OpenStudy (anonymous):

ehh i would put cannot combine because sr 10 is 2radical5 5sr5 is 5sr5 and 6sr5 is 6sr5 so is this a test or homework?

OpenStudy (anonymous):

Homework

OpenStudy (anonymous):

i would put cannot combine just my opinion